The mean weight of an adult is 62 kg with the variance of 144. if 195 adults are randomly selected what is the probability that the sample mean would be less than 59.5 kg found your answer to four decimal places
Answer #1

Solution :

Given that ,

mean = = 62

standard deviation = = 12

= / n = 12/ 195 = 0.8593

P( < 59.5) = P(( - ) / < (59.5 - 62) / 0.8593)

= P(z < -2.91)

= 0.0018

Probability = 0.0018